### v1

# Quantum Many-Body Scars for Arbitrary Integer Spin in $2+1$D Abelian Gauge Theories

Abstract
The existence of Quantum Many-Body Scars, high-energy eigenstates that evade the Eigenstate Thermalization Hypothesis, has been established across different quantum many-body systems, including gauge theories corresponding to spin-1/2 Quantum Link Models. We systematically identify scars for pure gauge theories with arbitrarily large integer spin $S$ in $2+1$D, concretely for Truncated Link Models, where the electric field is restricted to $12S+$ 1 states per link. Through an explicit analytic construction, we show that the presence of scars is widespread in $2+1$D gauge theories for arbitrary integer spin. We confirm these findings numerically for small truncated spin and $S=1$ Quantum Link Models. The proposed analytic construction establishes the presence of scars far beyond volumes and spins that can be probed with existing numerical methods.
Introduction.— Starting with an initial state and the microscopic Hamiltonian of an isolated system governing its dynamics, it might seem that the unitary evolution of the system would prevent thermalization. This puzzle is addressed by the Eigenstate Thermalization Hypothesis (ETH) [1, 2, 3, 4, 5]. According to the ETH, for non-integrable systems, the high-energy states and the observables of interest will eventually converge to a description in accordance with equilibrium statistical mechanics, achieving thermal equilibrium. In contrast to the ETH, many-body localized systems defy this trend due to their emergent integrability [6]. Systems exhibiting Quantum Many-Body Scars (QMBS) offer a contrasting example [7, 8, 9]. They also evade ETH, but only in an exponentially small fraction of states in the Hilbert space.
The effect of QMBS was first observed in a Rydberg-atom quantum simulator, marked by persistent revivals for particular initial states, in contrast to the vast majority of other high-energy initial states [10, 11]. In parallel with the exact construction of highly excited eigenstates in a non-integrable model [12, 13], the experiment has sparked intense research across different systems, where a weak breaking of the ETH takes place [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In this case, the presence of few anomalous eigenstates, can leave an imprint on thermalization. Among their special features is that they are characterized by atypically low entanglement entropy, compared with other states arbitrarily close in the spectrum. This is in stark contrast to what is expected from the ETH. In particular, QMBS may exhibit an area law. This means that the entanglement between a subsystem and its complement is proportional to the area of the boundary that divides them. In turn, highly excited states are expected to have an entanglement that grows linearly with the volume of the subsystem.
The interplay between gauge symmetry and scarring phenomena is not fully understood. Gauge theories naturally lead to constraints, and specific constraints can give rise to QMBS, as observed in the PXP model [10, 24, 25]. Gauge theories also feature local symmetries, which segment the Hamiltonian into distinct sectors that do not mix under time evolution. Supplied by a condition on physical states, this leads to local constraints in the form of Gauss’ law. Beyond the PXP model, QMBS have also been observed in other gauge theories, such as the Abelian case in the presence of matter in $1+1$D [26], pure gauge in $2+1$D [27, 28, 29] and non-Abelian in $2+1$D [30]. So far, the research on QMBS in Abelian pure gauge theories has exclusively focused on spin-$1/2$ Quantum Link Models (QLM) [27, 28, 29], which also exhibit other interesting phenomena such as crystalline confining phases [31]. This Letter demonstrates, for the first time, the existence of QMBS in link models with arbitrary integer spins.
Addressing questions related to real-time dynamics falls outside the capabilities of conventional lattice gauge theory methods, as Monte Carlo simulations face inefficiencies due to severe sign problems associated with real-time processes (see e.g. [32, 33]). Sign problems have motivated the development of the field of quantum simulations of gauge theories [34, 35, 36, 37, 38, 39]. Formulations like QLMs are essential for preserving gauge symmetry while ensuring a finite Hilbert space for a finite lattice volume. Recovering the theory in the continuum limit may take different routes [40, 41, 42, 43, 44, 45]. No matter the path, the road ahead is long until we can make decisive contributions to theories like Quantum Chromodynamics (QCD). Hence, it is imperative to pinpoint intriguing phenomena that can drive experiments along this path. The study of QMBS is a meritorious example as it is present in simple gauge theories and probes fundamental aspects of quantum many-body theory.
Demonstrating the presence of QMBS for large Hilbert spaces per gauge link, beyond spin-$1/2$, serves several fundamental purposes: it directly addresses how widespread QMBS are across many-body systems; it reveals novel mechanisms for their formation, illuminating the role played by gauge symmetry; and it helps to guide quantum simulation experiments toward interesting questions before the necessary complexity of QCD is achieved.
In this Letter, we demonstrate the extensive presence of QMBS across $2+1$ dimensional $U(1)$ gauge theories without matter. Concretely, we explicitly construct mid-spectrum states that satisfy area law entanglement for arbitrary integer spin, specifically for a simply truncated Hilbert space per link. We further verify the existence of these states by numerically determining the system’s eigenstates and calculating their entanglement and Shannon entropy for spin 1 and 2. Our findings unveil the presence of QMBS for single-leg ladders of the spin-1 QLM, which are more accessible to experiments than larger systems. Furthermore, our analytical approach allows us to identify QMBS for arbitrary integer spin and system size, circumventing the limitations of existing numerical methods when applied to large dimensional Hilbert spaces.
Link models in $2+1$D.— We consider models on square $L_{1}\times L_{2}$ lattices with bosonic degrees of freedom living on the links
$H=\sum_{n}\left(U_{n1}^{\dagger}U_{n+\hat{1}2}^{\dagger}U_{n2}U_{n+\hat{2}1}+\mathrm{h.c.}\right)+V,$ (1)
where the indices $n\equiv\left(n_{1},n_{2}\right)$ represent lattice sites, while the labels $i\in\left\{1,2\right\}$ represent the two directions. The first terms of the Hamiltonian are plaquette terms constructed by acting on each of the four links of a plaquette with operators represented by $U$. The term $V$, which we will call generically potential, will always be diagonal in the electric field $E_{ni}$. The commutation relations between these variables are $\left[E_{mi},U_{nj}\right]=U_{mi}\delta_{mn}\delta_{ij}$. Under these general conditions, the Hamiltonian has a set of local symmetries. Concretely, there is one generator of local gauge transformations $G_{n}$ per lattice site, which commutes with the Hamiltonian
$G_{n}=E_{n1}+E_{n2}-E_{n-\hat{1}1}-E_{n-\hat{2}2},\quad\left[H,G_{n}\right]=0.$ (2)
We will use the electric field basis, denoted by $\left|\varepsilon\right>$ for a single link. The generators $G_{n}$ are diagonal in this basis. The Hilbert space breaks into many different sectors. We will focus on the physical sector characterized by the Gauss’ law $G_{n}\left|\psi\right>=0$, for all sites, which acts as a constraint of the system.
The model also has two winding symmetries. Concretely, the observables $W_{1}=\sum_{n_{1}=0}^{L_{1}-1}E_{\left(n_{1},m\right)2}$ and $W_{2}=\sum_{n_{2}=0}^{L_{2}-1}E_{\left(m,n_{2}\right)1}$ commute with the Hamiltonian. In the physical sector, their eigenvalues are independent of $m$. We focus on the zero-winding sector $W_{i}\left|\psi\right>=0$ for both $i=1$ and $i=2$. This is the largest sector of the theory. There are further global symmetries that depend on the choice of the potential and are used to reduce the size of the Hilbert space in exact diagonalization calculations. They are described in the supplementary material [46]. We will be interested in two specific versions of this model that ensure that we will have a finite-dimensional Hilbert space per link.
In Quantum Link Models (QLM) [47, 48, 42] $U_{ni}/U_{ni}^{\dagger}$ are spin raising/lowering operators. $U_{ni}$ and $U^{\dagger}_{ni}$ do not commute and act according to $U\left|\varepsilon\right>\propto\sqrt{S\left(S+1\right)-\varepsilon\left(\varepsilon+1\right)}\ket{\varepsilon+1}$. The electric field operators $E$ correspond to the $z$ component spin operators.
In Truncated Link Models (TLM) [26] the electric field is simply truncated, meaning $U\left|S\right>=0$, $U^{\dagger}\left|-S\right>=0$ and $U\left|n\right>=\ket{n+1}$ otherwise. $U_{ni}$ and $U^{\dagger}_{ni}$ still do not commute, though the non-zero commutation relation is removed to the edge of the local Hilbert space (to the states |$\ket{-S}$ and |$\ket{S}$). With this construction, there will be $12S+$ 1 states per link. We generically refer to $S$ as the total spin, irrespective of the formulation we are using. The $S=1$ TLM is equivalent to the $S=1$ QLM.
For the Hamiltonian (1) we will also write
$H=K+V,\quad K=H^{+}+H^{-},\quad H^{\pm}=\sum_{n}H^{\pm}_{n},$ (3)
where $H^{-}_{n}=U_{n1}^{\dagger}U_{n+\hat{1}2}^{\dagger}U_{n2}U_{n+\hat{2}1}$ and $H^{+}=\left(H^{-}\right)^{\dagger}$ are kinetic operators. This notation follows the convention that we will adopt throughout the letter. “Plaquette at site $n$” refers to the plaquette for which $n$ is its lower left vertex. In order to simplify the formulas we will also adopt the convention $\left(H^{+}\right)^{-m}\equiv\left(H^{-}\right)^{m}$. This should not be confused with the inverse of $\left(H^{+}\right)^{m}$, since these operators are not invertible.
Zero-mode scars.— We refer to zero-modes as eigenstates of the plaquette term of the Hamiltonian with zero eigenvalues. In our notation, these states satisfy
$K\ket{\psi_{z}}=\sum_{n}(H^{+}_{n}+H^{-}_{n})\left|\psi_{z}\right>=0.$ (4)
Zero-modes can play a crucial role in forming quantum many-body scars. Systems with a spectral symmetry, together with point-group symmetries, can exhibit an exponential number of zero-modes [49]. Their existence follows from an index theorem, which we review and adapt to our models of interest in the supplementary material [46]. While one might expect these eigenstates to be thermal, it has been shown that low entangled states can be constructed within this subspace for specific cases. It was further conjectured that this is a generic property of local Hamiltonians with an exponential number of zero-modes [50]. Such a result was also observed in the case of gauge theories in the spin-$1/2$ QLM, even if it is possible to find scars that are not constructed exclusively from zero-modes [28, 29].
We show that it is also possible to construct such states for arbitrary spin truncated link models. The number of these low entropy states grows exponentially with volume. By choosing appropriately the potential term of the Hamiltonian, different linear combinations can be isolated as low entropy eigenstates. To establish these states as scars, one should prove that the model in question is non-integrable. Since we can construct these states for a wide region of parameter space, we expect this to be the case.
QMBS in TLM with arbitrary integer spin.— Next, we show that zero-modes with area-law entanglement entropy exist in truncated link models of all integer spins. We start by partitioning the 2D lattice into $2\times 1$ and $1\times 2$ tiles, as depicted on the left of Fig. 1. A given partition will be called tiling and will be denoted by $T$. It can be represented by a set of tuples $\left(n,n^{\prime}\right)$ where the two entries indicate the two plaquettes making up each tile. We then define the states
|$\ket{\psi_{s}^{(i,T)}}=$ $\frac{1}{\left(S+1\right)^{\left|T\right|/2}}\prod_{(n,n^{\prime})\in T}$
$\left(\sum_{k=0}^{S}(-1)^{k}(H^{+}_{n})^{i-S+k}(H^{+}_{n^{\prime}})^{i-k}\right)\ket{\mathbf{0}},$ (5)
where |$\ket{\mathbf{0}}$ is the state where all links have value zero, and $0\leq i\leq S$. The states |$\ket{\psi_{s}^{(i,T)}}$ have the property $(H^{-}_{m}+H^{+}_{m^{\prime}})\ket{\psi_{s}^{(i,T)}}=(H^{+}_{m}+H^{-}_{m^{\prime}})\ket{\psi_{s}^{(i,T)}}=0$ if $(m,m^{\prime})$ is a tile in $T$. Therefore,
$K\ket{\psi_{s}^{(i,T)}}=\sum_{(n,n^{\prime})\in T}(H^{+}_{n}+H^{-}_{n^{\prime}}+H^{-}_{n}+H^{+}_{n^{\prime}})\ket{\psi_{s}^{(i,T)}}=0,$ (6)
making them zero-modes.
Figure 1: Left: Example of a tiling $T$ of the 2D lattice, consisting of partitioning it into $2\times 1$ and $1\times 2$ tiles, to a total of 8 tiles. Right: Representation of a scar state based on an alternative tiling and in the dual representation. An independent sum over every $k_{i}$ with a weight of $\prod_{i}\left(-1\right)^{k_{i}}$ is assumed. See Eq. (19).
These states are zero-modes in all lattices where $L_{1}L_{2}$ is even. The number of possible tilings grows exponentially with $L_{1}L_{2}$ if both $L_{1}$ and $L_{2}$ are larger than 1.
By partitioning the systems into two regions, we see that the entanglement between them is generated by tiles that touch both regions resulting in an area law. The states |$\ket{\psi_{s}^{(i,T)}}$ are then mid-spectrum states with area law entanglement entropy, making them quantum scars.
To get a deeper understanding of the structure of these scars, and to connect them to lego scars that have been constructed in the $S=1/2$ QLM, it is useful to define the dual basis for integer spin. Height variables $h_{n}\in\mathbb{Z}$ live at the center of plaquettes. The values of the vertical links are given by the difference of the height variables of the plaquettes to their right and left, while the values of the horizontal links are given by the difference of the height variables of the plaquettes to their bottom and top, i.e. $E_{n1}=h_{n-\hat{2}}-h_{n}$ and $E_{n2}=h_{n}-h_{n-\hat{1}}$. Height variable configurations with $|h_{n}-h_{n^{\prime}}|\leq S$ for all neighboring plaquettes $n$, $n^{\prime}$ represent a valid spin-$S$ link configuration.
The height variable representation with periodic boundary conditions is not unique, since adding a constant to all height variables gives the same state. If there is an open boundary condition in at least one direction, we set all height variables at one open boundary to have the value of the neighboring boundary link. In our construction, we always choose the top boundary, which makes the representation unique. The kinetic operators $H^{+}_{n}$ and $H^{-}_{n}$, are raising/lowering operators for the height variable $h_{n}$.
In the dual representation, the state (19) in a $2\times 1$ plaquette system is
|$\ket{\psi_{i}}=\frac{1}{\sqrt{S+1}}\sum_{k=0}^{S}(-1)^{k}\ket{(i-S+k)\ (i-k)}$ (7)
where the state is labeled with its height variables |$\ket{h_{1}\ h_{2}}=(H^{+}_{1})^{h_{1}}(H^{+}_{2})^{h_{2}}\ket{\mathbf{0}}$. An illustration of this tile and the terms that cancel each other in the sum $K\ket{\psi_{i}}$ for $S=1$ can be found in Fig. 2.
Figure 2: Schematic depiction of the scar building block for $S=1$. The tiles depicted with solid lines make up the scar state. The operators $H^{\pm}_{n}$ raise or lower the height variables in plaquette $n$. They either annihilate the state or generate another state, depicted with dashed lines. This is canceled by applying the other type of operator on the second state.
All states that contribute in (19) have a height variable representation where there is one state out of the sum (7) on all tiles of the tiling $T$. The state sums over all $(S+1)^{\left|T\right|}$ combinations of these states. A pictorial representation of this tiling is shown on the right-hand side of Fig. 1. This tiling structure is similar to the lego and sub-lattice scars constructed for the $S=1/2$ case [28, 29], and allows for the construction of an exponential number of scars for arbitrary truncated link models. A formal definition of this tiling, in the form of a tiling product, can be found in the supplementary material [46]. There, we also describe a more general framework to construct more low-entropy zero-modes in TLMs.
Numerical results and discussion.— While these states are low entropy zero-modes, they are degenerate with all the other zero-modes and are not easily identifiable in the spectrum. In the following, we introduce a potential that has these states as eigenstates. By analyzing the Shannon entropy and the bipartite entanglement entropy of the eigenstates of the Hamiltonian, we verify the existence of these low entropy states numerically. See the supplementary material for a brief overview of the calculation of these quantities as used in this work [46].
From Eq. (7) we see that the sum of the height variables remains constant and equal to $(2i-S)$ over the entire sum. This does not mean that the sum over height variables is a valid potential because, in general, the height variables are not uniquely defined from the physical configuration. But they can be made unique for open boundary conditions, as described above. For example, in the case of a single-leg ladder, the sum of all height variables can be seen as the sum of the horizontal links of the top row. In general, for open boundary conditions, we write
$V=\lambda\sum_{n}h_{n},$ (8)
which will have the constructed scars as eigenstates with energy $E=\lambda\frac{L_{1}L_{2}}{2}(2i-S)$.
For $S=1$ the two zero-mode tiles are
[…]
which correspond to choosing $i=1$ and $i=0$ in (7). We first study a $S=1$ single-leg ladder with $L$ plaquettes. This is a chain of plaquettes with periodic boundary conditions in the direction of the chain, but not the direction perpendicular to it. For such a ladder, there are exactly two distinct tilings of $2\times 1$ tiles. They correspond to placing the lower left corner of the tiles in either even or odd sites. These two states, coming from two different tilings with the same $i$, are not orthogonal but generate a two-dimensional scarred Hilbert space. We then expect four scars. Two degenerate low entropy states with energy $L\lambda/2$ from the tilings of state |$\ket{\psi_{+}}$ ($i=1$) and two with energy $-L\lambda/2$ from tilings of state |$\ket{\psi_{-}}$ ($i=0$). This is exactly what is observed when performing exact diagonalization on small systems, the results of which are presented in Fig. 3. The four low-entropy states are clearly identifiable and each of them is composed of the expected states.
Figure 3: Left: Entanglement entropy of a $S=1$ ladder with $10\times 1$ plaquettes using the potential (8) with $\lambda=0.2$. Four mid-spectrum states with low entropy, shown in red, are visible. They are superpositions of the predicted states. For the remaining states, brighter colors indicate a higher density of states. Right: Entanglement entropy of $6\times 2$ $S=1$ plaquettes with periodic boundary conditions in the $x$ direction and $\lambda=0.2$ in the zero momentum sector. Many mid-spectrum states with low entropy are present. The towers marked in red contain the predicted states.
For systems with a larger vertical direction, there are many more ways to tile the plane with $2\times 1$ and $1\times 2$ tiles. Adding a second leg is enough to make the number of possible tilings exponentially large, all of which will mix in the spectrum. Using the sum of the height variables as the potential, we expect towers of low-entropy states at energy ±$\pm L_{1}L_{2}\lambda/2$. These two towers are observed, as depicted in Fig. 3. There are many other low-entropy states at other energies, which make use of vertical tiles. They entail a further structure, that allows the number of height variables with value $1$ not to be the same as the ones with value 00. While these are still scars built up from zero-modes, we do not expect this construction to be generalizable beyond $2\times L$ systems and will not attempt a detailed characterization here. For periodic lattices, we studied the $4\times 4$ system with a generalized version of the potential. There the expected scar towers are also present. Details for this case can be found in the supplementary material [46].
Instead of the designed potential that was used to isolate the constructed scars, one might want to use the standard electric field term
$V=g\sum_{n}E_{n}^{2}.$ (10)
This potential exhibits scars that are constructed from zero-modes and allows for the identification of new tiling patterns. We identify three low entropy states in a $4\times 4$ lattice, which turn out to be constructed from zero-modes (see Fig. 4). Two of these are generalizable to arbitrary $L_{1}\times L_{2}$ lattices if both $L_{i}$ are even. We characterize one of them here, the remaining two scars are described in the supplementary material [46].
The scar has the form (19), just as the other scars we derived. However, the two plaquettes that make up tiles do not share a link, but neighbor each other diagonally in a pattern depicted in Fig. 4. This is still a zero-mode since each plaquette has one neighbor with height ±$\pm 1$ in all configurations and may therefore never take height $\mp 1$ itself. The state is an eigenstate of (10) because the number of neighbors with height ±$\pm 1$ is always the same for both plaquettes of each tile. This means that in all configurations, half of the links have values ±$\pm 1$ and the others are 00. Establishing the presence of these scars beyond $S=1$ is beyond the scope of this work. Systems with larger spins are less accessible through ED beyond single-leg ladders and we did not observe any of these scars for $S\leq 2$ in this case.
Refer to caption         Refer to caption
Figure 4: Left: Shannon entropy of $4\times 4$ $S=1$ plaquettes with periodic boundary conditions using the $E^{2}$ potential (10) with $g=0.2$ in the zero momentum and $+1$ charge conjugation sectors. Three mid-spectrum states with low entropy, are visible and colored in red. Bright colors indicate a higher density of states. Right: Illustration of a $S=1$ zero-mode that is an eigenstate of $\sum_{n}E_{n}^{2}$. A sublattice structure is built by singlet states of two plaquettes connected diagonally. Plaquettes connected by red or green lines form tiles.
For $S=2$ we have studied the $6\times 1$ ladder in Fig. 5 with the height potential (8). In this case, we expect scars obtained from choosing $i=0$, $i=1$, and $i=2$ in (7). The $i=1$ scar has zero energy for any value of $\lambda$, while the other two will have $E_{2,1}=\pm 6\lambda$. The latter two are evident in the figure. The $i=1$ scar is not visible, since it is in superposition with a tower of zero-modes.
This demonstrates the existence of low entanglement entropy zero-modes in systems beyond $S=1$. For larger volume or higher spin systems, we have established theoretically the existence of such states in the earlier sections of this article.
Figure 5: Left: Entanglement entropy of a $S=2$ Ladder with 6 plaquettes with $\lambda=0.2$. Brighter colors indicate a higher density of states. Right: Shannon entropy of a $S=2$ Ladder with 6 plaquettes with $\lambda=0.2$. Four mid-spectrum states with low entropy are visible, colored in red. The remaining two predicted states are mixed with a tower of zero-modes at $E=0$.
Conclusions and outlook.— In this Letter, we demonstrate that QMBS exist across various 2D Abelian gauge theories. In the limit of zero coupling, a spectral symmetry emerges, leading to an exponentially large number of zero-modes. We show that it is possible to construct many states with area-law entanglement entropy, despite these states being in the middle of the spectrum. The number of these states grows exponentially with the volume, but they still comprise an exponentially small fraction of the Hilbert space, indicating weak ETH breaking.
We have analytically constructed QMBS for arbitrary truncated spins, revealing their existence beyond the reach of existing numerical methods, thereby marking a significant step towards understanding their role in the continuum limit, where the Hilbert space per link is not bounded. Our construction relies on the existence of numerous zero-modes, common in QLMs for arbitrary spin, though our method extends only to S=1 QLMs. This does not exclude that scars in higher spin QLMs exist but a different constructive approach might be needed. For $S=1$, some scars survive the presence of the standard $g\sum_{n}E^{2}_{n}$ potential. In a recent work [51], it has been argued that scars originally present on single-leg ladders $SU\left(2\right)$ gauge theories disappear once larger truncations of the Hilbert space are considered. Our results for $S=1$ show that these systems might be too simple to observe scars with the $E^{2}$ potential. An interesting generalization of our findings would involve extending the explicit QMBS construction to non-Abelian theories like $SU\left(2\right)$. From the experimental side, our theoretical description can guide future detection of QMBS and investigations of weak ergodicity breaking in systems inaccessible through classical simulations.
Acknowledgements.—We are grateful to Debasish Banerjee for insightful discussions. T.B. thanks the Galileo Galilei Institute for Theoretical Physics for the hospitality during the completion of part of this work. This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311. We acknowledge access to Piz Daint at the Swiss National Supercomputing Centre, Switzerland under the ETHZ’s share with the project ID eth8. Support from the Google Research Scholar Award in Quantum Computing and the Quantum Center at ETH Zurich is gratefully acknowledged.
Note.—During the final stages of our manuscript, we became aware of another work [52] on quantum many-body scarring in a 2+1D U(1) gauge theory with dynamical matter. This work will appear in the same arXiv listing as ours.